# Symmetry re-breaking in an effective theory of quantum coarsening

### v1

@@ Acknowledgements.
Appendix A Details of numerical simulations
@@
where $\nabla$ is the lattice derivative: $(\nabla\phi)_{(x,y)}\equiv(\phi_{(x+1,y)}-\phi_{(x,y)},\phi_{(x,y+1)}-\phi_{(x,y)})$ for $i\!=\!(x,y)$. The Hamilton equations of motion are
@@
Appendix B Gaussian decoupling of the $\phi^{4}$ theory
@@
where the prime on the sum indicates that the terms containing $\tilde{\phi}_{0}$ and $\tilde{\phi}_{0}^{2}$ were removed. Finally, we enforce gaussianity of the fluctuations by substituting $\tilde{\phi}_{k}^{4}$ with $3\expectationvalue*{\tilde{\phi}_{k}^{2}}\tilde{\phi}_{k}^{2}$ (Wick’s theorem), where the average $\expectationvalue*{\cdots}$ refers to an ensemble average over gaussian states:

### v2

@@ Acknowledgements.
Appendix A Algorithms for the numerical simulations
@@
where $\nabla$ is the lattice derivative: $(\nabla\phi)_{(x,y)}\equiv(\phi_{(x+1,y)}-\phi_{(x,y)},\phi_{(x,y+1)}-\phi_{(x,y)})$ for i=(x,y)i\!=\!(x,y). The Hamilton equations of motion are
@@
Appendix B Initial conditions for the numerical simulations
For the study of the areal velocity, we take as initial state for the dynamics a circular domain of radius r(t=0)=80r(t\!=\!0)\!=\!80 lattice sites, as in Fig. 2(a). The transverse field $g$ is fixed through the Hamiltonian parameters. To fix $\varepsilon$, we first let the outside state be the “down” ground state ($S_{i}^{x}\!=\!g/4$, $S_{i}^{z}\!=\!-\sqrt{1-g^{2}/16}$), and the one inside the disk to be the “up” ($S_{i}^{x}\!=\!g/4$, $S_{i}^{z}\!=\!+\sqrt{1-g^{2}/16}$). We then slowly heat up the configuration by coupling it to an effective bath (via Monte Carlo sampling), until it reaches the target excess energy density $\varepsilon$. Notice that the resulting configurations are not drawn from the Gibbs distribution, as we do not run the Monte Carlo algorithm for long enough. We finally time-evolve the resulting state.
For the study of the post-quench oscillations, we need to prepare the system at g=0g\!=\!0 and fixed excess energy density $\varepsilon_{0}\gtrsim 0$. Again, a nonzero excess energy density is obtained by warming up gradually the perfect ferromagnet $\vec{S}_{i}\!=\!+\hat{z}$ via Monte Carlo sampling.
Appendix C Mean-field frequency of oscillation
We show here how one can obtain the frequency $\omega(g)$ from Eq. (3) [main text]. The starting point is to realize that the mean-field, Lipkin-Meshkov-Glick Hamiltonian
$H_{\mathrm{LMG}}=-2(S^{z})^{2}-gS^{x}$ (10)
defines a classical Hamiltonian system. Canonically conjugated variables for the system can be taken to be $q,p$ such that
$S^{x}=p,\quad S^{y}=\sqrt{1-p^{2}}\sin q,\quad S^{z}=\sqrt{1-p^{2}}\cos q,$ (11)
with $\{q,p\}=1$ and $q\in[-\pi,\pi]$, $p\in[-1,1]$. In these new variables the Hamiltonian reads
$H_{\mathrm{LMG}}=-2(1-p^{2})\cos^{2}q-gp.$ (12)
One can now obtain the period of the trajectory at energy $E$ through the formula
$T(E)=\frac{d}{dE}\oint p\,dq.$ (13)
Notice that the energy must in the end be fixed to E=2E=-2, as the ferromagnetic initial condition corresponds to $q=0$, $p=0$. The curve $p(q;E)$ at energy $E$ is obtained by solving Eq. (12) wrt $p$:
$p_{\pm}(q;E)=\frac{g\pm\sqrt{g^{2}+8E\cos^{2}q+16\cos^{4}q}}{4\cos^{2}q}.$ (14)
The integral in Eq. (13) can be now expressed in terms of elliptic functions, but we found more convenient to just bring the derivative under the integral sign, and perform the integration numerically, finally obtaining $\omega=2\pi/T$ as displayed in Fig. 3(e) [main text].
Appendix D Gaussian decoupling of the $\phi^{4}$ theory
@@
[…]